A) A 20-year-old inherits $9,500 and puts it into a savings account earning 0.9% APR. How much can she expect this account to be worth when she turns 50, assuming a constant growth rate with regular monthly compounding?

B) A small family business will have a $1,000 tax liability at the end of one year. They would like to pay this liability with as little out-of-pocket money as possible. How much should they invest today in a Certificate of Deposit (CD) earning 3.5% APR yearly to accomplish this?
Answer #1

(A) Initial Inheritance = $ 9500, Tenure of Investment = 50 - 20 = 30 years or 360 months, APR = 0.9 % with monthly compounding

Monthly Rate = (0.9/12) = 0.075 %

Therefore, Accumulated Value at age 50 = 9500 x (1.00075)^(360) = $ 12443.4

(B) Tax Liability = $ 1000, Tenure = 1 year and CD Interest Rate = 3.5 % APR

Let the required current investment be $ P

Therefore, P x (1.035) = 1000

P = 1000 / (1.035) = $ 966.18